bill533:

Which answer best describes the complex zeros of the polynomial function? f(x)=x3+x2−8x−8 A. The function has one real zero and two nonreal zeros. The graph of the function intersects the x-axis at exactly one location. B. The function has one real zero and two nonreal zeros. The graph of the function intersects the x-axis at exactly two locations. C. The function has three real zeros. The graph of the function intersects the x-axis at exactly three locations. D. The function has two real zeros and one nonreal zero. The graph of the function intersects the x-axis at exactly one location.

bill533:

@silvernight269 @563blackghost

bill533:

@Shadow

Shadow:

\[f(x)=x^3+x^2−8x−8\] \[f(x) = x^2 (x + 1) -8(x +1)\] \[f(x) = (x^2 - 8)(x+1)\] \[0 = (x^2 - 8)(x+1)\] \[x^2 - 8 = 0\] \[x^2 = 8\] \[x = \pm \sqrt 8 = \pm 2 \sqrt 2\] \[x + 1 = 0\] \[x = -1\]

Shadow:

Does that work look familiar to what you have done in lessons?

bill533:

yes

Shadow:

Okay so then based on that, what do you think the answer is?

bill533:

the answer is A

Shadow:

The answer would be A if it were x^2 + 8, because that would make for: \[\pm 2i \sqrt 2\] Two imaginary numbers. But that is not the case.

bill533:

CORRECT

Shadow:

So if all the numbers are real, what would the answer be.

Shadow:

With: \[x = -1, 2 \sqrt 2, -2\sqrt 2\]

bill533:

A

Shadow:

The answer would be A if it were: \[x = -1, 2i \sqrt 2, -2i \sqrt 2\]

bill533:

oh D

Shadow:

"The function has two real zeros and one nonreal zero." But I said all the numbers are real. Lol stop guessing.

bill533:

no i was between both A or D because i knew B and C were out of the questions so it had to have been either A or D i thought it was D at first then thought A i should've followed my first mind

Shadow:

It's C cause all the zeroes are real...

bill533:

ok thank you on this shadow

bill533:

shadow could you help me on my new question please